2 times a number, increased by 4, is 12. Find the number
Answer by anjali_yamini(27)


Let :
THE NUMBER = N
SO, TWICE OF THE NUMBER = N+N= 2N

SO, TWICE THE NUMBER INCREASED BY 4 = 2N +4
SO, 2N +4 = 12
SO, 2N = 12-4
SO, 2N = 8
SO, N= 8/2
N = 4
CHECK :
2N + 4
= (2*4) +4
= 8 + 4
= 12